OpenStudy (anonymous):

The combustion of propane may be described as" C3H8 (g)+5O2(g)--> 3CO2(g)+4H2O(g) How many grams of O2 are needed to completely burn 36.2g of C2H8? G/O2

OpenStudy (aaronq):

find the amount of moles 36.2 g of C2H8 are. Use the stoichiometric coefficients to build a ratio, like this: e.g. for a general reaction: \(\color{red}{a}A + \color{blue}{b}B\) \(\rightleftharpoons\) \( \color{green}{c}C\) where upper case are the species (A,B,C), and lower case (a,b,c) are the coefficients (in red), \(\dfrac{n_A}{\color{red}{a}}=\dfrac{n_B}{\color{blue}{b}}=\dfrac{n_C}{\color{green}{c}}\) from here, you can solve for moles then convert to grams. To convert mass to moles, use the relationship: \(n=\dfrac{m}{M}\) where, M=molar mass, m=mass, and n= moles.

OpenStudy (anonymous):

I don't quite understand the second part

OpenStudy (aaronq):

post what you've done and i'll help you correct it